OpenStudy (anonymous):

graph the linear equation. -3x+9y=18 ?

OpenStudy (anonymous):

@FilthyMcNasty ??

OpenStudy (anonymous):

-3x+9y=18 -3(0)+9y=18 9y=18 y=2 -3x+9(0)=18 -3x=18 x=-6

OpenStudy (anonymous):